Find the sum of 1st 12 term of the series 3,7,11,15,19
Answer by mananth(16946)

3,7,11,15
is an Arithmetic sequence
a=3 and d=4
n=12
Sn= n/2[2a+(n-1)d]
S12 =12/2[2*3+(12-1)*4]
=6[6+44]
=6*50
=300
S12 = 300